OpenStudy (anonymous):

Give the foci of x^2/9+y^2/4=1 (sq.rt of____,____),(-sq rt. of___,____)

OpenStudy (anonymous):

the foci is obtained using the value of c in c^2 = a^2 - b^2, where a=3, b=2. once you find the value of c, move c units up and down from the center to get your foci.

OpenStudy (anonymous):

can you give the value of c?

OpenStudy (anonymous):

c= sq. root of 13

OpenStudy (anonymous):

subtract.. don't add...

OpenStudy (anonymous):

c^2 = a^2 - b^2 = 3^2 - 2^2 = 9 - 4 = 5 so c = sqrt5

OpenStudy (anonymous):

plus/minus sqrt5

OpenStudy (anonymous):

also... sorry 'bout the mistake up there... the center is (0, 0) and you go left/right of the center, NOT up/down of the center.

OpenStudy (anonymous):

okay thanks, is it the same for a hyperbola? Do I add or subtract?